# Spin-Refined Partition Functions and $\mathcal{CRT}$ Black Holes

### v1

@@ Abstract
We investigate spin-refined partition functions in AdS/CFT using Euclidean gravitational path integrals. We construct phase diagrams for $Z_{X}=\text{Tr}\big{(}e^{-\beta H}X\big{)}$ in various dimensions and for different choices of discrete isometry $X$, discovering rich structures at finite temperature. When $X$ is a reflection, $Z_{X}$ counts the difference between the number of even- and odd-spin microstates. The high-temperature regime is universally dominated by $\mathcal{CRT}$-twisted black holes in any dimension, and in odd spacetime dimensions we examine whether complex rotating black hole solutions could contribute to spin-refined observables. We also analyze the microcanonical ensemble. There the leading contribution almost always come from rotating black holes, showing that the two ensembles are not equivalent.
@@ 1 Introduction
In recent years there has been a surge of activity in trying to understand black hole microstates using the gravitational path integral Gibbons:1976ue ; Cabo-Bizet:2018ehj ; Iliesiu:2022kny ; Iliesiu:2022onk ; LopesCardoso:2022hvc ; H:2023qko ; Sen:2023dps ; Anupam:2023yns , especially in AdS/CFT Maldacena:1997re ; Witten:1998zw . The main interest lies in computing the thermal partition function $Z=\textrm{Tr}\big{(}e^{-\beta H}\big{)}$, and also the density of states, but these are not the only probes of the statistics one can employ. The most famous alternative is the Witten index Witten:1982df $\textrm{Tr}\big{(}e^{-\beta H}(-1)^{\mathsf{F}}\big{)}$ and its generalizations (see for example Sen:2012hv ). The index counts the relative number of bosonic to fermionic states by tracing over fermionic states with the opposite sign. An even more general setting was considered in Benjamin:2019stq , where the twisted partition function
$\textrm{Tr}\big{(}e^{-\beta H+2\pi i\alpha J}\big{)}$ (1)
was considered in 2d CFTs, with the special attention given to the case $\alpha\in\mathbb{Q}$. Observables of this form, which we will call spin-refined partition functions, have the potential to give more fine-grained insight into the microstate structure of gravitational theories. For similar results in the presence of a finite-group symmetry, see Harlow:2021trr .
@@
$Z_{\vec{\alpha}}=\textrm{Tr}e^{-\beta H+\vec{\alpha}\cdot\vec{J}},$ (2)
where $\vec{J}=(J_{1},...,J_{n})$ is the angular momentum taking values in the Cartan subalgebra of the rotation algebra $\mathfrak{so}(D-2)$, and $\vec{\alpha}$ is a set of fixed numbers. Most often we will assume that all of them are rational multiples of $2\pi i$. We will also consider an additional insertion of the reflection operator $R$. To get a better feeling of the quantities we will be studying, let us start by providing a microscopic interpretation of
$Z_{R}(\beta)=\textrm{Tr}e^{-\beta H+i\pi J}$ (3)
in 2d CFTs. If the theory living on the boundary has reflection symmetry, then for every eigenvector |$\ket{E,j}$ of the hamiltonian $H$ and angular momentum $J$, there exists also an eigenvector |$\ket{E,-j}$. If $j\in\frac{1}{2}+\mathbb{Z}$, the contributions from $j$ and $-j$ differ by a phase that cancels them out exactly. Thus, $Z_{R}$ is a purely bosonic quantity. [^1] Note that this does not mean that the presence of fermions in the theory or lack thereof is meaningless. Rather, only bosonic states contribute, but they may still be built out of fermions. For bosons with even $j$, $e^{i\pi J}$ is positive, and for bosons with odd $j$ it is negative. Thus we may rewrite $Z_{R}$ as
@@
Notice that on $S^{1}$, the antipodal map is the same as the rotation $\phi\mapsto\phi+\pi$. Thus, in this case we have $Z_{R}(\beta)=\textrm{Tr}e^{-\beta H}R$. We will in fact take this as the definition of $Z_{R}$ in any dimension. We may still interpret it as (4), where now the labels “even” and “odd” refer to the properties of states under $R$. If the boundary theory has $R$ as a symmetry, it will be a bosonic quantity in any dimension.
@@
1.1 A review of the $\mathcal{CRT}$-twisted black hole
The authors of Harlow:2023hjb provided an example of a universal saddle contributing to $Z_{R}$ that we will now review. Let us start with the non-rotating BTZ black hole with the inverse temperature $\beta$. In Kruskal-Szekeres coordinates $(X^{\pm},\phi)$, the metric reads
@@ 1.1 A review of the $\mathcal{CRT}$-twisted black hole
For future reference, note that on $S^{1}$, the rotation $\phi\mapsto\phi+\pi$ is an antipodal map. It is a property of odd-dimensional spheres that the antipodal map is just a rotation.
Having established that the $\mathcal{CRT}$-twisted black hole is a saddle for the gravitational path integral, we should also establish what boundary conditions it satisfies. The time circle gets cut in half, so its inverse temperature is $\frac{\beta}{2}$. Moreover, the identification introduces a twist in a spatial circle. It was argued in Harlow:2023hjb that this saddle contributes to $\textrm{Tr}\big{(}e^{-\frac{\beta}{2}H+i\pi J}\big{)}$. [^3] It may be worth emphasizing that this is not an index, since $(-1)^{\mathsf{F}}=e^{2\pi iJ}$. Notice that this corresponds to the modular parameter $\tau=\frac{1}{2}+i\frac{\beta}{4\pi}$. The partition function at this value of the modular parameter can be rewritten using modular invariance as
@@
A few observations regarding that problem are in place. One may be worried that a large split between even and odd $j$ could lead to problems with the semiclassical limit of the bulk theory. After all, we do not expect that the scattering process of a photon on the Kerr background should depend significantly on whether the spin of the black hole is $j$ or $j+\hbar$. However, while the $Z_{R}(\beta)$ may be large, the relative change in the density of states
$\frac{Z_{R}(\beta)}{Z(\beta)}$ (13)
is exponentially small as $G_{\text{N}}\to 0$. Thus, the semiclassical physics should be agnostic to that split.
@@ 2.1 Thermodynamics of rotating systems
One traditionally starts the discussion of thermodynamics with the entropy $S$ as a function of extensive parameters [^5] Although these parameters are not necessarily extensive in black hole thermodynamics, the same ideas apply.—in our case, the energy $E$ and the angular momentum $J$ [^6] In general spacetime dimension $D$, angular momentum takes values in the Cartan subalgebra of $O(D-2)$, and black hole thermodynamics is modified accordingly The first law of black hole thermodynamics then reads then
@@
The trace is taken over states with energy [^7] To be more precise, we trace over states lying within a small window of energies centered at $E$. $E$, such that for each state $\beta\Omega$ is fixed to a prescribed value. We have, then,
@@
Therefore $S_{\text{rot}}$ is in a natural way a function of $E$ and $\beta\Omega=:\alpha$ (which is consistent with the fact that in the trace expression we keep that combination fixed).
@@
$S_{CRT}(\beta)=\frac{1}{2}S_{\textrm{Schwarzschild}}(2\beta).$ (23)
@@
$E_{\mathcal{CRT}}(\beta)=E_{\textrm{Schwarzschild}}(2\beta).$ (24)
@@
$S_{\mathcal{CRT}}(E)=\frac{1}{2}S_{\textrm{Schwarzschild}}(E).$ (25)
It follows in particular that in any dimension, the $\mathcal{CRT}$ black hole contribution to $\textrm{Tr}_{E}R$ is equal to $\sqrt{\textrm{Tr}_{E}1}$. As we will see below, it is far from being the dominating contribution though (at least at large energies).
@@ 2.2 An example: BTZ black holes
In the fixed temperature and angular velocity ensemble, which we will call the canonical ensemble from now on, the free energy of rotating BTZ reads
$F_{rot}=\frac{\pi^{2}c}{3\beta^{2}(\Omega^{2}-1)}.$ (26)
From this we may easily calculate the entropy [^8] One may wonder why we bother with the thermodynamical derivations of (for example) entropy, since this is proportional to the horizon’s area and thus easily obtained from the geometry. The reason is that we treat it as a warm-up before the case of a general $SL(2,\mathbb{Z})$ black hole, where these expressions may be less familiar.
$S=-\left(\frac{\partial F_{rot}}{\partial T}\right)_{\Omega}=\frac{2\pi^{2}c}{3\beta(1-\Omega^{2})},$ (27)
@@
$J=-\left(\frac{\partial F_{rot}}{\partial\Omega}\right)_{\beta}=\frac{2\pi^{2}\Omega c}{3\beta^{2}(1-\Omega^{2})^{2}},$ (28)
@@
$S_{rot}=S+\alpha J$ (30)
treated as a function of $E$ and $\alpha$. The analytical result is not very illuminating. What actually matters to us is the behavior of $S_{rot}$ as $E\to\infty$. We find that for $\alpha^{2}E\gg c$, we have
$S_{rot,\pm\pm}(E,\alpha)=\pm\alpha E\pm\frac{1}{\sqrt{2}}S_{0}(E)+O(1),$ (31)
@@
Let $\alpha=i\tilde{\alpha}$, so that $\tilde{\alpha}\in\mathbb{R}$. The quantity we are actually interested in is $\textrm{Tr}_{E}\big{(}e^{i\tilde{\alpha}J}\big{)}$, and this is determined by $e^{S_{rot}}$ for the branch of (31) with the largest real part. The first term in $S_{rot,\pm\pm}$ is purely imaginary and describes oscillations in the density of states, while the second term is the leading real part. The solutions $S_{rot,\pm,+}$ are dominant, while $S_{rot,\pm-}$ are exponentially suppressed. Thus, at large energies we may write
@@
First, however, note that $\tilde{\alpha}$ is defined only mod $2\pi$999or $4\pi$, if the theory includes fermions. We will discuss the inclusion of fermions below. It follows (for example, from modular invariance) that we should include all possible discrete values of $\tilde{\alpha}$. If $\tilde{\alpha}$ happens to be 00 (meaning the angular velocity is zero), then saddles with $\tilde{\alpha}+2\pi n$ $(n\neq 0)$ will be subleading. However, if $\tilde{\alpha}\notin 2\pi\mathbb{Z}$, then all possible values of $\tilde{\alpha}$ contribute at the same order: the difference between them is $O\big{(}\frac{1}{\sqrt{E}}\big{)}$.
@@ 2.3 The $SL(2,\mathbb{Z})$ black holes
$Z_{C,D}(\tau)=Z_{AdS}\left(\frac{A\tau+B}{C\tau+D}\right)=\exp\left(\frac{\pi c}{6}\Im\left(\frac{A\tau+B}{C\tau+D}\right)\right)=\exp\left(\frac{\pi^{2}c}{3}\frac{\beta}{(\tilde{\alpha}C+2\pi D)^{2}+C^{2}\beta^{2}}\right).$ (34)
@@
$F_{rot}=-\frac{\pi^{2}c}{3}\frac{1}{(\tilde{\alpha}C+2\pi D)^{2}+C^{2}\beta^{2}}.$ (35)
Notice that when $\tilde{\alpha}\in\pi\mathbb{Q}$, there is a special modular transformation such that $\tilde{\alpha}C+2\pi D=0$. We will come back to this shortly. To keep the ‘real’ notation from the previous section, write $i\tilde{\alpha}=\alpha=\beta\Omega$. We may use standard themodynamical relations to get
$J=-\left(\frac{\partial F_{rot}}{\partial\Omega}\right)_{\beta}=\frac{\pi^{2}c}{3}\frac{2i\beta C(2\pi D-iC\beta\Omega)}{\big{(}\beta^{2}C^{2}+(2\pi D-iC\beta\Omega)^{2}\big{)}^{2}},$ (36a)
$S=-\left(\frac{\partial F_{rot}}{\partial T}\right)_{\Omega}=-\frac{\pi^{2}c}{3}\frac{2\beta^{2}C\left(\beta C\left(\Omega^{2}-1\right)+2i\pi D\Omega\right)}{\left(\beta^{2}C^{2}\left(\Omega^{2}-1\right)+4i\pi\beta CD\Omega-4\pi^{2}D^{2}\right)^{2}},$ (36b)
@@
$E=F_{rot}+TS+\Omega J=\frac{\pi^{2}c}{3}\frac{C^{2}\left(\alpha^{2}+\beta^{2}\right)+4i\pi\alpha CD-4\pi^{2}D^{2}}{\left(C^{2}\left(\beta^{2}-\alpha^{2}\right)-4i\pi\alpha CD+4\pi^{2}D^{2}\right)^{2}}.$ (36c)
As before, the last equation (where we have put $\alpha=\beta\Omega$) can be solved for $\beta(E,\alpha)$ (there are again four solutions $\beta_{\pm\pm}$) and then plugged into $S_{rot}=S+\alpha J$, treated as a function of $E$ and $\alpha$. The behavior of $S_{rot}$ as $E\to\infty$, for generic values of $\alpha,C,D$ is
$S_{rot,\pm\pm}(E,\alpha)=\pm i\left(\frac{2D\pi-iC\alpha}{C}\right)E\pm\frac{1}{\sqrt{2}C}S_{0}(E)+O(1).$ (37)
Setting $\alpha=i\tilde{\alpha}$, we see that once again the first term above is purely imaginary and describes oscillations. The second term is the leading real part, and as before $S_{rot,\pm+}$ is dominant while $S_{rot,\pm-}$ is suppressed. Notice, however, that for $C>1$, $S_{rot,\pm\pm}(E,\alpha)$ is always subleading with respect to $C=1$ (which corresponds to the BTZ black hole). Thus, at large energies the microcanonical ensemble is always dominated by the BTZ black hole.
@@
$S_{rot}(E)=\frac{2\pi\sqrt{cE}}{\sqrt{3}C}=\frac{S_{0}}{C}+O(1).$ (38)
Even this expression is subleading with respect to the BTZ black hole. However, is has one important feature: it is purely real. Thus, its contribution to the density of states does not oscillate. It follows that if we consider the microcanonical ensemble with sufficiently wide energy windows, this is the contribution that will survive while all others may cancel out. We will see that this is exactly what happens in the thermal partition function.

### v2

@@ Abstract
We investigate spin-refined partition functions in AdS/CFT using Euclidean gravitational path integrals. We construct phase diagrams for $Z_{X}=\text{Tr}\big{(}e^{-\beta H}X\big{)}$ in various dimensions and for different choices of discrete isometry $X$, discovering rich structures at finite temperature. When $X$ is a reflection, $Z_{X}$ counts the difference between the number of even- and odd-spin microstates. The high-temperature regime is universally dominated by $\mathcal{CRT}$-twisted black holes in any dimension, and in odd spacetime dimensions we examine whether complex rotating black hole solutions can contribute to spin-refined observables or potentially dominate at finite temperature. We also analyze the microcanonical ensemble. There the leading contribution almost always comes from rotating black holes, showing that the two ensembles are not necessarily equivalent.
@@ 1 Introduction
In recent years there has been a surge of activity in trying to understand black hole microstates using gravitational path integrals Gibbons:1976ue ; Cabo-Bizet:2018ehj ; Iliesiu:2022kny ; Iliesiu:2022onk ; LopesCardoso:2022hvc ; H:2023qko ; Sen:2023dps ; Anupam:2023yns , especially in AdS/CFT Maldacena:1997re ; Witten:1998zw . The main interest lies in computing the thermal partition function $Z=\textrm{Tr}\big{(}e^{-\beta H}\big{)}$ and the density of states, but these are not the only probes of statistics that one can employ. The most famous alternative is the Witten index $\textrm{Tr}\big{(}e^{-\beta H}(-1)^{\mathsf{F}}\big{)}$ Witten:1982df , along with its generalizations (see for example Sen:2012hv ). The index counts the difference between the number of bosonic and fermionic states in the theory by tracing over fermions with the opposite sign. An even more general quantity was considered in Benjamin:2019stq , where the twisted partition function $\textrm{Tr}\big{(}e^{-\beta H+\alpha J}\big{)}$ was studied in 2d conformal field theories (CFTs), with special attention given to the case $\alpha\in 2\pi i\mathbb{Q}$. Observables of this form, which we will call spin-refined partition functions, have the potential to give more fine-grained insight into the microstate structure of gravitational theories. For similar results in the presence of global and finite-group symmetries, see Pal:2020wwd ; Harlow:2021trr .
@@
$Z_{\vec{\alpha}}(\beta)=\textrm{Tr}\left(e^{-\beta H+\vec{\alpha}\cdot\vec{J}}\right),$ (1)
where $\vec{J}=(J_{1},...,J_{n})$ is the angular momentum, taking values in the Cartan subalgebra of the rotation algebra $\mathfrak{so}(D-2)$, and $\vec{\alpha}$ is a set of fixed numbers. Most often we will assume that all of them are rational multiples of $2\pi i$. To get a better feeling for the quantities we will be studying, let us start by providing a microscopic interpretation of
$Z_{R}(\beta)=\textrm{Tr}\left(e^{-\beta H+\pi iJ}\right)$ (2)
in 2d CFTs. If the theory living on the boundary has reflection symmetry, then for every eigenvector |$\ket{E,j}$ of the Hamiltonian $H$ and angular momentum $J$, there exists also an eigenvector |$\ket{E,-j}$. If $j\in\frac{1}{2}+\mathbb{Z}$, the contributions from $j$ and $-j$ differ by a phase that cancels them out exactly. Thus, $Z_{R}$ is a purely bosonic quantity. [^1] Note that this does not mean that the presence of fermions in the theory or lack thereof is meaningless. Rather, only bosonic states contribute, but these states may still be built out of fermions. For bosons with even $j$, $e^{\pi iJ}$ is positive, and for bosons with odd $j$ it is negative. Thus we may rewrite $Z_{R}$ as
@@
Notice that on $S^{1}$, the antipodal map is the same as the rotation $\phi\mapsto\phi+\pi$. More generally, it is a property of odd-dimensional spheres that the antipodal map is just a rotation by $\pi$. Thus, in this case we have $Z_{R}(\beta)=\textrm{Tr}\big{(}e^{-\beta H}R\big{)}$, where $R$ is the reflection operator. [^2] We use $R$ to denote an insertion that reflects of all of the coordinates. Thus $R=\cal{RT}$, where $\cal{R}$ and $\cal{T}$ are the discrete symmetries corresponding to spatial reflection and (Euclidean) time reversal. We will in fact take this as the definition of $Z_{R}$ in any dimension. We may still interpret it as (3), where now the labels “even” and “odd” refer to the properties of states under $R$. If the boundary theory has $R$ as a symmetry, then $Z_{R}$ will be a bosonic quantity in any dimension.
@@
Let us pause to emphasize that in our usage, the operator $R$ changes the signs of all the coordinates. This may be confusing, because in Euclidean signature the parity operator flips the signs of all spatial coordinates, while the reflection operator flips only one of them. But parity and reflection are actually the same in odd bulk spacetime dimensions, up to a rotation. (For example, in $\text{AdS}_{3}/\text{CFT}_{2}$ they are identical.) Our focus in this work will be on examples in odd spacetime dimensions, so we will not be very careful with the distinction. [^4] The reflection of only one spatial coordinate has determinant $-1$ in any dimension, so Euclidean saddles for that insertion can be non-orientable. Such objects have been studied in $\text{AdS}_{3}/\text{CFT}_{2}$ Maloney:2016gsg ; Wei:2024zez , but these are not the saddles we are looking for. We thank Zixia Wei for clarifying comments on this point. In practice, we take the definition of $R$ to be such that it implements (3).
1.1 Review of the $\mathcal{CRT}$-twisted black hole
The authors of Harlow:2023hjb provided an example of a universal saddle contributing to $Z_{R}$ that we will now review. Let us start with the non-rotating BTZ black hole with inverse temperature $\beta$. In Kruskal-Szekeres coordinates $(X^{\pm},\phi)$, the metric reads
@@ 1.1 Review of the $\mathcal{CRT}$-twisted black hole
Having established that the $\mathcal{CRT}$-twisted black hole is a saddle for the gravitational path integral, we should also establish what boundary conditions it satisfies. The time circle gets cut in half, so its inverse temperature is $\frac{\beta}{2}$. Moreover, the identification introduces a twist in a spatial circle. It was argued in Harlow:2023hjb that this saddle contributes to $\textrm{Tr}\big{(}e^{-\frac{\beta}{2}H+\pi iJ}\big{)}$. [^5] It may be worth emphasizing that this is not an index, since $(-1)^{\mathsf{F}}=e^{2\pi iJ}$. Notice that this corresponds to the modular parameter $\tau=\frac{1}{2}+i\frac{\beta}{4\pi}$. The partition function at this value of the modular parameter can be rewritten using modular invariance as
@@
A few observations regarding that problem are in place. One may be worried that a large split between even and odd $j$ could lead to problems with the semiclassical limit of the bulk theory. After all, we do not expect that the scattering process of a photon on the Kerr background should depend significantly on whether the spin of the black hole is $j$ or $j+\hbar$. However, while $Z_{R}(\beta)$ may be large, the relative change in the density of states $Z_{R}(\beta)/Z(\beta)$ remains exponentially small as $G_{\text{N}}\to 0$. Thus, the semiclassical physics should be agnostic to the even-odd split.
@@ 2.1 Thermodynamics of rotating systems
One traditionally starts discussions of thermodynamics with the entropy $S$ as a function of extensive parameters [^7] Although these parameters are not necessarily extensive in black hole thermodynamics, the same ideas apply.—in our case, the energy $E$ and the angular momentum $J$ [^8] In general spacetime dimension $D$, angular momentum takes values in the Cartan subalgebra of the rotation algebra $\mathfrak{so}(D-2)$, and black hole thermodynamics is modified accordingly. The first law of black hole thermodynamics then reads
@@
The trace is taken over states with fixed energy $E$, such that for each state $\beta\Omega$ is fixed to a prescribed value. [^9] The notation $\mathrm{Tr}_{E}$ means that we trace over states in a small window of energies $(E,E+\delta E)$. This typically gives microcanonical quantities which agree with Legendre transforms of appropriate canonical quantities. However, oscillatory quantities like $e^{\pi iJ}$ are not positive-definite, so changing the window slightly can alter the results. However, with a sufficiently wide window, there is a natural sense in which such quantities oscillate around a central value. This is the perspective taken in Chen:2023mbc , which also tried to quantify the window-dependence of oscillatory quantities by studying wormhole contributions to their connected correlator. We have, then,
@@
Therefore $S_{\text{rot}}$ is in a natural way a function of $E$ and $\beta\Omega=:\alpha$, which is consistent with the fact that in the trace expression we keep that combination fixed.
@@
$S_{CRT}(\beta)=\frac{1}{2}S_{\textrm{SAdS}}(2\beta).$ (21)
@@
$E_{\mathcal{CRT}}(\beta)=E_{\textrm{SAdS}}(2\beta).$ (22)
@@
$S_{\mathcal{CRT}}(E)=\frac{1}{2}S_{\textrm{SAdS}}(E).$ (23)
It follows in particular that in any dimension, the $\mathcal{CRT}$ black hole contribution to $\textrm{Tr}_{E}R$ is equal to $\sqrt{\textrm{Tr}_{E}1}$. However, as we will see below, this is far from being the dominant contribution, at least at large energies.
@@ 2.2 An example: BTZ black holes
In the ensemble of fixed temperature and angular velocity, which we will call the canonical ensemble from now on, the free energy of the rotating BTZ black hole reads
$F_{\text{rot}}=\frac{\pi^{2}c}{3\beta^{2}(\Omega^{2}-1)}.$ (24)
From this we may easily calculate the entropy [^10] One may wonder why we bother deriving (for example) the entropy, since this is proportional to the horizon’s area and thus easily obtained from the geometry. The reason is that we treat it as a warm-up before the case of a general $SL(2,\mathbb{Z})$ black hole, where these expressions may be less familiar.
$S=-\left(\frac{\partial F_{\text{rot}}}{\partial T}\right)_{\Omega}=\frac{2\pi^{2}c}{3\beta(1-\Omega^{2})},$ (25)
@@
$J=-\left(\frac{\partial F_{\text{rot}}}{\partial\Omega}\right)_{\beta}=\frac{2\pi^{2}\Omega c}{3\beta^{2}(1-\Omega^{2})^{2}},$ (26)
@@
$S_{\text{rot}}=S+\alpha J$ (28)
treated as a function of $E$ and $\alpha$. The analytical result is not very illuminating. What actually matters to us is the behavior of $S_{\text{rot}}$ as $E\to\infty$. We find that for $\alpha^{2}E\gg c$, we have
$S_{\text{rot},\pm\pm}(E,\alpha)=\pm\alpha E\pm\frac{1}{\sqrt{2}}S_{0}(E)+O(1),$ (29)
@@
Let $\alpha=i\tilde{\alpha}$, so that $\tilde{\alpha}\in\mathbb{R}$. The quantity we are actually interested in is $\textrm{Tr}_{E}\big{(}e^{i\tilde{\alpha}J}\big{)}$, and this is determined by $e^{S_{\text{rot}}}$ for the branch of (29) with the largest real part. The first term in $S_{\text{rot},\pm\pm}$ is purely imaginary and describes oscillations in the density of states, while the second term is the leading real part. The solutions $S_{\text{rot},\pm,+}$ are dominant, while $S_{\text{rot},\pm-}$ are exponentially suppressed. Thus, at large energies we may write
@@
First, however, note that $\tilde{\alpha}$ is defined only mod $2\pi$111111or $4\pi$, if the theory includes fermions. We will discuss the inclusion of fermions below. It follows (for example, from modular invariance) that we should include all possible discrete values of $\tilde{\alpha}$. If $\tilde{\alpha}$ happens to be 00 (meaning the angular velocity is zero), then saddles with $\tilde{\alpha}+2\pi n$ $(n\neq 0)$ will be subleading. However, if $\tilde{\alpha}\notin 2\pi\mathbb{Z}$, then all possible values of $\tilde{\alpha}$ contribute at the same order: the difference between them is $O\big{(}\frac{1}{\sqrt{E}}\big{)}$.
@@ 2.3 The $SL(2,\mathbb{Z})$ black holes
$Z_{C,D}(\tau)=Z_{\text{AdS}}\left(\frac{A\tau+B}{C\tau+D}\right)=\exp\left(\frac{\pi c}{6}\Im\left(\frac{A\tau+B}{C\tau+D}\right)\right)=\exp\left(\frac{\pi^{2}c}{3}\frac{\beta}{(\tilde{\alpha}C+2\pi D)^{2}+C^{2}\beta^{2}}\right).$ (32)
@@
$F_{\text{rot}}=-\frac{\pi^{2}c}{3}\frac{1}{(\tilde{\alpha}C+2\pi D)^{2}+C^{2}\beta^{2}}.$ (33)
Notice that when $\tilde{\alpha}\in\pi\mathbb{Q}$, there is a special modular transformation such that $\tilde{\alpha}C+2\pi D=0$. We will come back to this shortly. To keep the ‘real’ notation from the previous section, write $i\tilde{\alpha}=\alpha=\beta\Omega$. We may use standard thermodynamical relations to get
$J=-\left(\frac{\partial F_{\text{rot}}}{\partial\Omega}\right)_{\beta}=\frac{\pi^{2}c}{3}\frac{2i\beta C(2\pi D-iC\beta\Omega)}{\big{(}\beta^{2}C^{2}+(2\pi D-iC\beta\Omega)^{2}\big{)}^{2}},$ (34a)
$S=-\left(\frac{\partial F_{\text{rot}}}{\partial T}\right)_{\Omega}=-\frac{\pi^{2}c}{3}\frac{2\beta^{2}C\left(\beta C\left(\Omega^{2}-1\right)+2\pi iD\Omega\right)}{\left(\beta^{2}C^{2}\left(\Omega^{2}-1\right)+4\pi i\beta CD\Omega-4\pi^{2}D^{2}\right)^{2}},$ (34b)
@@
$E=F_{\text{rot}}+TS+\Omega J=\frac{\pi^{2}c}{3}\frac{C^{2}\left(\alpha^{2}+\beta^{2}\right)+4\pi i\alpha CD-4\pi^{2}D^{2}}{\left(C^{2}\left(\beta^{2}-\alpha^{2}\right)-4\pi i\alpha CD+4\pi^{2}D^{2}\right)^{2}}.$ (34c)
As before, the last equation (where we have put $\alpha=\beta\Omega$) can be solved for $\beta(E,\alpha)$ (there are again four solutions $\beta_{\pm\pm}$) and then plugged into $S_{\text{rot}}=S+\alpha J$, treated as a function of $E$ and $\alpha$. The behavior of $S_{\text{rot}}$ as $E\to\infty$, for generic values of $\alpha,C,D$ is
$S_{\text{rot},\pm\pm}(E,\alpha)=\pm i\left(\frac{2D\pi-iC\alpha}{C}\right)E\pm\frac{1}{\sqrt{2}C}S_{0}(E)+O(1).$ (35)
Setting $\alpha=i\tilde{\alpha}$, we see that once again the first term above is purely imaginary and describes oscillations. The second term is the leading real part, and as before $S_{\text{rot},\pm+}$ is dominant while $S_{\text{rot},\pm-}$ is suppressed. Notice, however, that for $C>1$, $S_{\text{rot},\pm\pm}(E,\alpha)$ is always subleading with respect to $C=1$ (which corresponds to the BTZ black hole). Thus, at large energies the microcanonical ensemble is always dominated by the BTZ black hole.
@@
$S_{\text{rot}}(E)=\frac{2\pi\sqrt{cE}}{\sqrt{3}C}=\frac{S_{0}}{C}+O(1).$ (36)
Even this expression is subleading with respect to the BTZ black hole. However, is has one important feature: it is purely real. Thus, its contribution to the density of states does not oscillate. It follows that if we consider the microcanonical ensemble with sufficiently wide energy windows, this is the contribution that will survive while all others may cancel out. We will see that this is exactly what happens in the thermal partition function. We are thus ready to answer the question posed in the Introduction: are there systematically more even than odd spin states? Looking at Eq. (30), we see that this is not the case at leading order: indeed, $\textrm{Tr}_{E}\left(e^{\pi iJ}\right)$ is a highly oscillatory function. However, the important point is that there is also a subleading contribution, given by (36), which does not oscillate. This is the reason why there is a split between odd and even spins in the canonical ensemble.